# Timer: Transformers for Time Series Analysis at Scale

### v1

@@ 4.5 Analysis
Flexible sequence length
The decoder-only architecture provides additional flexibility to accommodate series of different lengths. For example, Timer can handle variable lookback lengths less than the context length, since it has been supervisedly trained on each context length during pre-training as Equation 4, thereby achieving enhanced performance by the increasing lookback length in Figure 11. It can also generate arbitrary tokens as the prediction by autoregressively sliding. By contrast, realizing this by encoder-only models with the rigid context length would necessitate rolling forecasting, leading to significant error accumulation, which is illustrated at the bottom of Figure 11.
Figure 11: Performance of Timer on variable context length.
Domain transfer
To investigate the effectiveness of domain partitioning of UTSD, we examined the results of adopting different domains of UTSD as the source and adapting to different target datasets to establish in-domain and out-of-domain transfer. The results in Table 2 indicate that in-domain transfer can further enhance the downstream performance. Additionally, as the number of downstream data samples increases, the relative improvement of pre-training will gradually diminish, and even lead to negative transfer in some out-of-domain scenarios. It provides a promising direction to develop domain-specific pre-trained models.
5 Conclusion
Time series analysis scenarios increasingly underscore the demand for large time series models. In this paper, we construct time series corpora from diverse domains, standardize them into a unified sequence format, and utilize a generative objective to pre-train a decoder-only Transformer, presenting a scalable, task-agnostic Timer for time series analysis. Experimentally, we evaluate the generality of Timer in forecasting, imputation, and anomaly detection, yielding remarkable downstream results and notable pre-training benefits in data-scarce scenarios. Further analysis validates the model scalability and effectiveness of architecture, and flexibility to tackle variable series lengths. In the future, we will continue to scale up the datasets and Timer, as well as explore advanced capabilities of LTSMs, such as the zero-shot generalization and in-context learning ability, to address a wider range of time series analysis tasks.
Appendix A Unified Time Series Dataset
A.1 Datasets Details
Unified Time Series Dataset (UTSD) is meticulously assembled from a blend of publicly accessible online data repositories and empirical data derived from real-world machine operations. To enhance data integrity, missing values are systematically addressed using linear interpolation techniques. We follow the unified data storage format used in Autoformer (Wu et al., 2021). For each univariate, multivariate and irregular-sampled time series, we align the time series with timestamps and store them in one single CSV file. One dataset may composed of multiple related time series. UTSD encompasses a collection of 29 individual datasets as listed in Table 3, intricately representative of a wide range of domains.
All datasets are classified into seven distinct domains by their source: Energy, Environment, Health, Internet of Things (IoT), Nature, Transportation, and Web. Datasets of UTSD exhibit diverse sampling frequencies, ranging from macro intervals such as yearly and quarterly to more fine-grained intervals like hourly and minutely. Notably, several datasets can demonstrate exceptionally high-frequency sampling rates, such as the MotorImagery dataset, which operates at a millisecond frequency.
In the pursuit of advanced data analysis, we have also analyzed the stationarity manifested as ADF test statistics (Elliott et al., 1996) and forecastability (Goerg, 2013). The rigorous methodologies and intricate details are elaborated in Section A.2.
Table 3: Dataset detailed descriptions. Time Point denotes the total number of time points aggregating from all variates if multivariate. File Size denotes the storage that the CSV file of the dataset occupies. Freq. denotes the sampling interval of time points, where “-” indicates no timestamp or irregular interval. ADF. denotes the Augmented Dickey-Fuller test statistics of the dataset. Forecast. denotes the forecastability of the dataset. Source denotes the original paper or resource of the dataset.
Domain Dataset Time Points File Size Freq. ADF. Forecast. Source
Energy
London Smart Meters
166.50M 4120M Hourly -13.158 0.173
Godahewa et al. (2021)
Wind Farms
7.40M 179M 4 sec -29.174 0.811
Godahewa et al. (2021)
Aus. Electricity Demand
1.16M 35M 30 min -27.554 0.730
Godahewa et al. (2021)
Environment
AustraliaRainfall
11.54M 54M Hourly -150.10 0.458
Tan et al. (2021)
BeijingPM25Quality
3.66M 26M Hourly -31.415 0.404
Tan et al. (2021)
BenzeneConcentration
16.34M 206M Hourly -65.187 0.526
Tan et al. (2021)
Health
MotorImagery
72.58M 514M 0.001 sec -3.132 0.449
Dau et al. (2019)
SelfRegulationSCP1
3.02M 18M 0.004 sec -3.191 0.504
Dau et al. (2019)
SelfRegulationSCP2
3.06M 18M 0.004 sec -2.715 0.481
Dau et al. (2019)
AtrialFibrillation
0.04M 1M 0.008 sec -7.061 0.167
Dau et al. (2019)
PigArtPressure
0.62M 7M - -7.649 0.739
Dau et al. (2019)
PigCVP
0.62M 7M - -4.855 0.577
Dau et al. (2019)
IEEEPPG
15.48M 136M 0.008 sec -7.725 0.380
Tan et al. (2021)
BIDMC32HR
63.59M 651M - -14.135 0.523
Tan et al. (2021)
TDBrain
72.30M 1333M 0.002 sec -3.167 0.967
Wang et al. (2023)
IoT
SensorData
165.4M 2067M 0.02 sec -15.892 0.917
Real-world machine logs
Nature
Phoneme
2.16M 25M - -8.506 0.243
Dau et al. (2019)
EigenWorms
27.95M 252M - -12.201 0.393
Dau et al. (2019)
ERA5 Surface
58.44M 574M 3 h -28.263 0.493
Muñoz-Sabater et al. (2021)
ERA5 Pressure
116.88M 1083M 3h -22.001 0.853
Muñoz-Sabater et al. (2021)
Temperature Rain
23.25M 109M Daily -10.952 0.133
Godahewa et al. (2021)
StarLightCurves
9.46M 109M - -1.891 0.555
Dau et al. (2019)
Saugeen River Flow
0.02M 1M Daily -19.305 0.300
Godahewa et al. (2021)
KDD Cup 2018
2.94M 67M Hourly -10.107 0.362
Godahewa et al. (2021)
US Births
0.00M 1M Daily -3.352 0.675
Godahewa et al. (2021)
Sunspot
0.07M 2M Daily -7.866 0.287
Godahewa et al. (2021)
Worms
0.23M 4M 0.033 sec -3.851 0.395
Dau et al. (2019)
Transport
Pedestrian Counts
3.13M 72M Hourly -23.462 0.297
Godahewa et al. (2021)
Web
Web Traffic
116.49M 388M Daily -8.272 0.299
Godahewa et al. (2021)
A.2 Statistics
Stationarity
Forecastability
Forecastability is calculated by subtracting the entropy of the series Fourier decomposition adopted from Goerg (2013), where a higher forecastability value indicates superior predictability. Just as with the assessment of stationarity, when considering a dataset composed of multiple time series of varying lengths, it is essential to adjust the measure of forecastability to account for these differences. Therefore, we extend the concept of forecastability to a weighted version, analogous to the length-weighted ADF method, to finely tune the predictability assessment to the individual characteristics of each series. The weighted forecastability for a dataset can be formulated as follows:
A.3 UTSD Composition Analysis
Dataset Complexity
We conduct a comprehensive analysis of individual datasets to obtain the stationarity and forecastability measures and construct the UTSD hierarchically regarding these indicators. Concretely, based on the ${\operatorname{ADF-Statistic}}$ of each dataset, we categorized the predictive difficulty of the datasets into three distinct levels: Easy, Medium, and Hard. The classification criteria are as follows:
  • Easy: ${\operatorname{ADF-Statistic}}<-15.00;$
  • Medium: $-15.00\leq{\operatorname{ADF-Statistic}}<-5.00;$
  • Hard: $-5.00\leq{\operatorname{ADF-Statistic}}.$
Figure 12: Dataset complexity in each hierarchy of UTSD.
Pattern Diversity
Each individual dataset within the UTSD collection demonstrates unique patterns, highlighting the importance of maintaining pattern diversity. We build the UTSD dataset in a top-down order, ensuring that each hierarchy within UTSD comprehensively represents all individual datasets and contains as many patterns as possible. As shown in Figure 13, we select several representative datasets for visualization analysis:
  • AtrialFibrillation: this dataset showcases a fluctuating trend with minimal seasonality. This pattern is indicative of irregular heart rhythm characteristics, typical in medical recordings related to cardiac health. Such fluctuations, lacking a clear seasonal pattern, are crucial for understanding the unpredictable nature of atrial fibrillation.
  • PigArtPressure: this dataset reveals a fluctuating trend interspersed with notable seasonality. This pattern is representative of the physiological variations in blood pressure that can occur due to environmental or internal factors. The presence of both fluctuating trends and seasonality in this dataset underscores the complex nature of biological data.
  • US Births: this dataset distinctly exhibits a clear trend alongside pronounced seasonality. This pattern is characteristic of demographic data, where trends and seasonal variations can reflect socio-cultural factors and environmental influences. The consistent trend and seasonality in birth rates provide insights into population dynamics and reproductive behaviors.
To avoid selecting trivial temporal variations and provide a comprehensive representation of the varied patterns inherent in the individual datasets, we employ a downsampling technique for individual datasets. For those with a larger number of variates, we selectively choose representative variates that best encapsulate the distinct patterns of the respective individual datasets. Similarly, for individual datasets with considerable temporal length, we extract the representative period. This methodical selection process ensures that the integrity and distinctive characteristics of each individual dataset are preserved, thereby maintaining the diversity of patterns across the hierarchical structure of the UTSD dataset.

### v2

@@ 4.4 Scalability
4.5 Model Analysis
Backbone for LTSM
To make sure the evaluation is comparable on different backbones, we maintain the same model configuration, including the model dimension and layer number, and pre-train these backbones on our UTSD-4G respectively. We set the token length $S=96$ and the context length as $672$ for Timer. For other non-autoregressive backbones, we pre-train them by direct multi-step forecasting in the $672$-pred-$96$ setting. The loss curves of training and validation are calculated as the MSE of the same set of model outputs (length-$96$ time series). As illustrated in Figure 12, Transformer exhibits excellent scalable ability as the backbone for LTSMs, whereas MLP-based and CNN-based architectures may encounter the bottleneck in accommodating diverse time series data.
Decoder-only v.s. Encoder-only
While a smaller training loss is achieved by the encoder-only Transformer in Figure 12, the progress of large language models indicates that decoder-only models may possess stronger generalization capabilities in downstream adaptation (Wang et al., 2022a; Dai et al., 2022), which is the essential purpose of LTSMs. Therefore, we proceed to compare their forecasting performance under varying degrees of data scarcity.
We elaborately evaluate two architectures on six benchmarks in Table 1. In the case of training from scratch (Pre-trained = None), the encoder-only Transformer will achieve better performance if the training samples are insufficient (Target = 1%). Instead, the decoder-only architecture will demonstrate improved performance when more training samples are provided in the end-to-end scenarios. After pre-training on UTSD-12G (Pre-trained = 12G), Timer as the decoder-only Transformer achieves the best performance in most downstream scenarios, indicating better generalization than the encoder-only pre-trained model. The observations are consistent with several findings in large language models and elucidate why the encoder-only structure has become prevalent in the field of time series. Existing benchmarks can still be small and the encoder-only model can overfit in end-to-end scenarios. Meanwhile, the decoder-only Transformer, which excels at generalizing on different domains, is a promising choice for developing large time series models.
Figure 13: Performance of one Timer for all lookback lengths.
Figure 14: Zero-shot evaluation on LTSMs. The top three models for each dataset are highlighted on the leaderboard. Average Rank of each model is calculated on the benchmarks in which the model has participated. Detailed results are provided in Table 18.
Flexible sequence length
Typically, current deep forecasting models are trained on specific lookback and forecast lengths, limiting their versatility. Instead, the decoder-only architecture can provide the flexibility to address different sequence lengths. For instance, one Timer is applicable on different lookback lengths because of token-wise supervision outlined in Equation 4. In addition to the feasibility, it achieves enhanced performance by increasing the lookback length in Figure 13. As for the forecast length, increasing works (Liu et al., 2024) bring the revival of autoregression (iterative multi-step prediction), enabling the generation of future predictions with arbitrary lengths. We explore this paradigm by rolling one model for all forecast lengths in Figure 15, where the decoder-only Times exhibits smaller error accumulation, thereby achieving better performance.
Figure 15: Performance of Timer/PatchTST for all forecast lengths. We conduct rolling forecasting on a single $672$-pred-$96$ model.
4.6 Evaluation of Large Time Series Models
There is a growing surge in the development of large models in the field of time series (Garza & Mergenthaler-Canseco, 2023; Das et al., 2023b; Woo et al., 2024; Ansari et al., 2024; Goswami et al., 2024). One particularly fascinating direction of research is focused on zero-shot forecasting (ZSF), which has the potential to renovate the conventional practice of training small models or fine-tuning language models for each specific scenario. Zero-shot generalization represents a sophisticated capability of large models, necessitating substantial model capacity and pre-training on extensive datasets. Consequently, we are actively expanding our dataset by incorporating the latest data infrastructure (Woo et al., 2024) in this field to pre-train Timer on ever larger scales (1B/16B/28B). Given the significant value to researchers and practitioners, we extensively evaluate concurrent large models and establish the first zero-shot forecasting benchmark of LTSMs as detailed in Appendix B.2.
Quality assessments
Our evaluation assesses the quality of LTSMs in Table 9, including (1) fundamental attributes such as pre-training scale, parameters; (2) abilities such as applicable tasks, context length, etc. Current LTSMs essentially build upon Transformer, with a significantly smaller number of parameters compared to LLMs. There is still potential to support more tasks and longer contexts.
Quantitative evaluations
We apply official checkpoints on seven datasets that do not appear during pre-training. The performance is fairly evaluated using MSE by predicting future $96$ points of all windows in each dataset. Figure 14 presents the result and rank of each model, where the top-ranked LTSMs are Timer, Moiria (Woo et al., 2024), and TimesFM (Das et al., 2023b). However, the positive correlation between performance and pre-training scale remains relatively weak, highlighting the significance of high-quality data and synchronized scaling of data and model size.
5 Conclusion and Future Work
Real-world time series analysis is increasingly underscoring the demand for large time series models (LTSM). In this paper, we release a time series dataset with 1 billion time points, propose a unified sequence format to address the heterogeneity of multivariate time series, and develop a generative pre-trained Transformer as a generalizable, scalable, task-general LTSM. Empirically, we evaluate our model in forecasting, imputation, and anomaly detection, yielding state-of-the-art performance and notable pre-training benefits in the data-scarce scenario. Further analysis validates the model scalability, explores the architecture for LTSMs, and highlights the versatility of our autoregressive generation. By performing zero-shot forecasting on available large models, we conduct the initial quantitative assessments among LTSMs. Quality evaluations unveil crucial pathways for future development, including better zero-shot generalization and facilitating probabilistic and long-context forecasting.
Impact Statement
This paper aims to advance the development of large models for time series. In this work, we release a high-quality and unified time series dataset for scalable pre-training, which can serve as a foundation for pre-training and establishing new benchmarks. The outcome large model demonstrates notable effectiveness of generalization, versatility across various tasks, and scalability to refine performance, offering valuable insights for future investigations and application values for practitioners. Our paper mainly focuses on scientific research and has no obvious negative social impact.
Acknowledgements
This work was supported by the National Natural Science Foundation of China (62022050 and U2342217), the BNRist Innovation Fund (BNR2024RC01010), and the National Engineering Research Center for Big Data Software.
Appendix A Unified Time Series Dataset
A.1 Datasets Details
Unified Time Series Dataset (UTSD) is meticulously assembled from a blend of publicly accessible online data repositories and empirical data derived from real-world machine operations. To enhance data integrity, missing values are systematically addressed using linear interpolation techniques. We follow the unified data storage format (parquet) used in  (Woo et al., 2024). For each univariate, multivariate, or irregular-sampled time series, we store them with timestamps and other meta-information in one directory using ARROW format. One dataset may composed of multiple related time series. We continue to expand the UTSD to include data from public datasets such as LOSTA [^1] https://huggingface.co/datasets/Salesforce/lotsa_data for zero-shot forecasting. UTSD encompasses 29 individual datasets as listed with the asterisk mark in Table LABEL:tab:dataset_detailed_descriptions, intricately representative of a wide range of domains.
All datasets can be classified into ten distinct domains by their source: Energy, Environment, Health, Internet of Things (IoT), Nature, Transport, Web, CloudOps, Finance, and Multiple Sources (Misc.), where the first seven domains originally come from our curated UTSD. The datasets exhibit diverse sampling frequencies, ranging from macro intervals such as yearly and quarterly to more fine-grained intervals like hourly and minutely. Notably, several datasets can demonstrate exceptionally high-frequency sampling rates, such as the MotorImagery dataset, which operates at a millisecond frequency.
In the pursuit of advanced data analysis, we have also analyzed the stationarity manifested as ADF test statistics (Elliott et al., 1996) and forecastability (Goerg, 2013). The rigorous methodologies and intricate details are elaborated in Section A.2. We utilize these statistical indicators to filter four high-quality subsets of UTSD, namely UTSD-1G, UTSD-2G, UTSD-4G, and UTSD-12G. As we expand the dataset, we continuously analyze statistical indicators and employ various methodologies to ensure the selection of high-quality datasets. LOTSA has not been sorted in this hierarchy due to its immensity.
A.2 Statistics
Stationarity
Forecastability
Forecastability is calculated by subtracting the entropy of the series Fourier decomposition adopted from Goerg (2013), where a higher forecastability value indicates superior predictability. Just as with the assessment of stationarity, when considering a dataset composed of multiple time series of varying lengths, it is essential to adjust the measure of forecastability to account for these differences. Therefore, we extend the concept of forecastability to a weighted version, analogous to the length-weighted ADF method, to finely tune the predictability assessment to the characteristics of each series. The weighted forecastability for a dataset can be formulated as follows:
{ThreePartTable}{TableNotes}
The asterisk marks the dataset that originally belongs to UTSD.
Table 2: Dataset detailed descriptions. Time Points denotes the total number of time points aggregating from all variates if multivariate. File Size denotes the storage that the ARROW format of the dataset occupies. Freq. denotes the sampling interval of time points, where “-” indicates no timestamp or irregular interval. ADF. denotes the Augmented Dickey-Fuller test statistics of the dataset. Forecast. denotes the forecastability of the dataset. Source denotes the original paper or resource of the dataset.
Domain Dataset Time Points File Size Freq. ADF. Forecast. Source
Energy
London Smart Meters*
166.50M 636M Hourly -13.158 0.173
Godahewa et al. (2021)
Wind Farms*
7.40M 29M 4 sec -29.174 0.811
Godahewa et al. (2021)
Aus. Electricity Demand*
1.16M 5M 30 min -27.554 0.730
Godahewa et al. (2021)
BDG-2 Panther
0.92M 4M Hourly -6.593 0.479
Emami et al. (2023)
BDG-2 Fox
2.32M 9M Hourly -9.191 0.469
Emami et al. (2023)
BDG-2 Rat
4.73M 19M Hourly -6.868 0.456
Emami et al. (2023)
BDG-2 Bear
1.48M 6M Hourly -11.742 0.471
Emami et al. (2023)
Low Carbon London
9.54M 37M Hourly -12.366 0.134
Emami et al. (2023)
SMART
0.10M 1M Hourly -10.755 0.143
Emami et al. (2023)
IDEAL
1.26M 5M Hourly -11.223 0.106
Emami et al. (2023)
Sceaux
0.03M 1M Hourly -14.172 0.143
Emami et al. (2023)
Borealis
0.08M 1M Hourly -6.612 0.160
Emami et al. (2023)
Buildings900K
15852.22M 60102M Hourly -8.412 0.357
Emami et al. (2023)
Covid19 Energy
0.03M 1M Hourly -13.768 0.698
Wang et al. (2023b)
GEF12
1.58M 6M Hourly -9.576 0.566
Wang et al. (2023b)
GEF14
0.02M 1M Hourly -9.372 0.628
Wang et al. (2023b)
GEF17
0.28M 1M Hourly -5.976 0.599
Wang et al. (2023b)
PDB
0.04M 1M Hourly -6.453 0.622
Wang et al. (2023b)
Spanish
0.07M 1M Hourly -13.217 0.770
Wang et al. (2023b)
ELF
0.02M 1M Hourly -13.607 0.770
Wang et al. (2023b)
KDD Cup 2022
4.73M 181M Hourly -17.017 0.225
Zhou et al. (2022)
Residential Load Power
437.98M 1671M Minutely -37.979 0.264
Bergmeir et al. (2023)
Residential PV Power
373.37M 1435M Minutely -31.389 0.421
Bergmeir et al. (2023)
Environment
AustraliaRainfall*
11.54M 45M Hourly -150.10 0.458
Tan et al. (2021)
BeijingPM25Quality*
3.66M 14M Hourly -31.415 0.404
Tan et al. (2021)
BenzeneConcentration*
16.34M 63M Hourly -65.187 0.526
Tan et al. (2021)
China Air Quality*
34.29M 132M Hourly -12.602 0.529
Zheng et al. (2015)
Beijing Air Quality*
4.62M 18M Hourly -15.758 0.332
Chen (2019)
Health
MotorImagery*
72.58M 279M 0.001 sec -3.132 0.449
Dau et al. (2019)
SelfRegulationSCP1*
3.02M 12M 0.004 sec -3.191 0.504
Dau et al. (2019)
SelfRegulationSCP2*
3.06M 12M 0.004 sec -2.715 0.481
Dau et al. (2019)
AtrialFibrillation*
0.04M 1M 0.008 sec -7.061 0.167
Dau et al. (2019)
PigArtPressure*
0.62M 3M - -7.649 0.739
Dau et al. (2019)
PigCVP*
0.62M 3M - -4.855 0.577
Dau et al. (2019)
IEEEPPG*
15.48M 61M 0.008 sec -7.725 0.380
Tan et al. (2021)
BIDMC32HR*
63.59M 244M - -14.135 0.523
Tan et al. (2021)
TDBrain*
72.30M 283M 0.002 sec -3.167 0.967
Wang et al. (2023a)
CDC Fluview ILINet
0.28M 2M Weekly -4.381 0.307
CDC (2017)
CDC Fluview WHO NREVSS
0.14M 1M Weekly -7.928 0.233
CDC (2017)
Project Tycho
1.35M 5M Weekly -8.167 0.111
van Panhuis et al. (2018)
IoT
SensorData*
165.4M 631M 0.02 sec -15.892 0.917
Real-world machine logs
Nature
Phoneme*
2.16M 9M - -8.506 0.243
Dau et al. (2019)
EigenWorms*
27.95M 107M - -12.201 0.393
Dau et al. (2019)
ERA5*
96458.81M 368610M Hourly -7.970 0.581
Nguyen et al. (2024)
CMIP6
104593.00M 399069M 6 H -7.960 0.573
Nguyen et al. (2024)
Temperature Rain*
23.25M 93M Daily -10.952 0.133
Godahewa et al. (2021)
StarLightCurves*
9.46M 37M - -1.891 0.555
Dau et al. (2019)
Saugeen River Flow*
0.02M 1M Daily -19.305 0.300
Godahewa et al. (2021)
KDD Cup 2018*
2.94M 12M Hourly -10.107 0.362
Godahewa et al. (2021)
US Births*
0.00M 1M Daily -3.352 0.675
Godahewa et al. (2021)
Sunspot*
0.07M 1M Daily -7.866 0.287
Godahewa et al. (2021)
Worms
0.23M 1M 0.033 sec -3.851 0.395
Dau et al. (2019)
Subseasonal
56.79M 217M Daily -12.391 0.414
Mouatadid et al. (2024)
Subseasonal Precipitation
9.76M 38M Daily -13.567 0.276
Mouatadid et al. (2024)
Transport
Pedestrian Counts*
3.13M 12M Hourly -23.462 0.297
Godahewa et al. (2021)
PEMS 03
9.38M 36M 5 min -19.051 0.411
Jiang et al. (2023)
PEMS 04
15.65M 60M 5 min -15.192 0.494
Jiang et al. (2023)
PEMS 07
24.92M 96M 5 min -20.603 0.466
Jiang et al. (2023)
PEMS 08
9.11M 35M 5 min -14.918 0.551
Jiang et al. (2023)
PEMS Bay
16.94M 65M 5 min -12.770 0.704
Jiang et al. (2023)
Los-Loop
7.09M 28M 5 min -16.014 0.657
Jiang et al. (2023)
Transport
Loop Seattle
33.95M 130M 5 min -32.209 0.535
Jiang et al. (2023)
SZ-Taxi
0.46M 2M 15 min -5.900 0.217
Jiang et al. (2023)
Beijing Subway
0.87M 22M 30 min -8.571 0.219
Jiang et al. (2023)
SHMetroy
5.07M 20M 15 min -17.014 0.222
Jiang et al. (2023)
HZMetro
0.38M 2M 15 min -11.254 0.232
Jiang et al. (2023)
Q-Traffic
264.39M 1011M 15 min -15.761 0.490
Jiang et al. (2023)
Taxi
55.00M 212M 30 min -8.302 0.146
Alexandrov et al. (2020)
Uber TLC Daily
0.05M 1M Daily -1.778 0.285
Alexandrov et al. (2020)
Uber TLC Hourly
1.13M 5M Hourly -9.022 0.124
Alexandrov et al. (2020)
LargeST
4452.20M 16988M 5 min -38.020 0.436
Liu et al. (2023a)
Web
Web Traffic*
116.49M 462M Daily -8.272 0.299
Godahewa et al. (2021)
Wiki-Rolling
40.62M 157M Daily -5.524 0.242
Alexandrov et al. (2020)
CloudOps
Alibaba Cluster Trace 2018
190.39M 2909M 5 min -5.303 0.668
Woo et al. (2023)
Azure VM Traces 2017
885.52M 10140M 5 min -11.482 0.290
Woo et al. (2023)
Borg Cluster Data 2011
1073.89M 14362M 5 min -8.975 0.505
Woo et al. (2023)
Sales
M5
58.33M 224M Daily -6.985 0.247
Alexandrov et al. (2020)
Favorita Sales
139.18M 535M Daily -6.441 0.097
Kaggle
Favorita Transactions
0.08M 1M Daily -5.481 0.362
Kaggle
Restaurant
0.29M 2M Daily -4.650 0.126
Kaggle
Hierarchical Sales
0.21M 1M Daily -8.704 0.078
Mancuso et al. (2021)
Finance
GoDaddy
0.26M 2M Monthly -1.539 0.784
Kaggle
Bitcoin*
0.07M 1M Daily -2.493 0.398
Godahewa et al. (2021)
Misc.
M1 Yearly
0.00M 1M Yearly -0.791 0.473
Godahewa et al. (2021)
M1 Quarterly
0.01M 1M Quarterly -0.175 0.572
Godahewa et al. (2021)
M1 Monthly
0.04M 1M Monthly -1.299 0.588
Godahewa et al. (2021)
M3 Yearly
0.02M 1M Yearly -0.850 0.524
Godahewa et al. (2021)
M3 Quarterly
0.04M 1M Quarterly -0.897 0.624
Godahewa et al. (2021)
M3 Monthly
0.1M 1M Monthly -1.954 0.635
Godahewa et al. (2021)
M3 Other
0.01M 1M - -0.568 0.801
Godahewa et al. (2021)
M4 Yearly
0.84M 4M Yearly -0.036 0.533
Godahewa et al. (2021)
M4 Quarterly
2.214M 10M Quarterly -0.745 0.696
Godahewa et al. (2021)
M4 Monthly
10.38M 41M Monthly -1.358 0.665
Godahewa et al. (2021)
Misc.
M4 Weekly
0.37M 2M Weekly -0.533 0.644
Godahewa et al. (2021)
M4 Daily
9.96M 39M Daily -1.332 0.841
Godahewa et al. (2021)
M4 Hourly
0.35M 2M Hourly -2.073 0.532
Godahewa et al. (2021)
[…]
Figure 16: Dataset complexity in each hierarchy of UTSD.
A.3 UTSD Composition Analysis
Dataset Complexity
We conduct a comprehensive analysis of individual datasets to obtain the stationarity and forecastability measures and construct the UTSD hierarchically regarding these indicators. Code for calculating the statistics is provided in the repository of UTSD. Consequently, based on the ${\operatorname{ADF-Statistic}}$ of each dataset, we categorized the predictive difficulty of the datasets into three levels: Easy, Medium, and Hard. The criteria are listed as follows:
  • Easy: ${\operatorname{ADF-Statistic}}<-15.00;$
  • Medium: $-15.00\leq{\operatorname{ADF-Statistic}}<-5.00;$
  • Hard: $-5.00\leq{\operatorname{ADF-Statistic}}.$
Pattern Diversity
Each dataset within the UTSD collection demonstrates unique patterns, highlighting the importance of maintaining pattern diversity. We build the UTSD dataset in a top-down manner, ensuring that each hierarchy within UTSD comprehensively represents all individual datasets and contains as many patterns as possible. As shown in Figure 17, we select several representative datasets for visualization analysis:
  • AtrialFibrillation: The dataset showcases a fluctuating trend with minimal seasonality. This pattern is an indicator of irregular heart rhythm characteristics, typical in medical recordings related to cardiac health. Such fluctuations, lacking a clear seasonal pattern, are crucial for understanding the unpredictable nature of atrial fibrillation.
  • PigArtPressure: The dataset reveals a fluctuating trend interspersed with notable seasonality. This pattern is representative of the physiological variations in blood pressure that can occur due to environmental or internal factors. The presence of both fluctuating trends and seasonality in this dataset underscores the complex nature of biological data.
  • US Births: The dataset distinctly exhibits a clear trend alongside pronounced seasonality. This pattern is characteristic of demographic data, where trends and seasonal variations can reflect socio-cultural factors and environmental influences. The consistent trend and seasonality in birth rates provide insights into population dynamics and reproductive behaviors.
To avoid selecting trivial temporal variations and provide a comprehensive representation of the varied patterns inherent in the individual datasets, we employ a downsampling technique for individual datasets. For those with a larger number of variates, we selectively choose representative variates that best encapsulate the distinct patterns of respective datasets. Similarly, for datasets with considerable temporal length, we resample them by the representative period. This methodical selection process ensures that the integrity and distinctive characteristics of each dataset are preserved, thereby maintaining the diversity of patterns across the hierarchical structure of the UTSD dataset.
A.4 Experiments
Forecasting benchmarks
In the field of time series forecasting, several classical datasets such as ETT (Zhou et al., 2021), ECL (Wu et al., 2021), Traffic (Wu et al., 2021) and Weather (Wu et al., 2021) have become widely recognized benchmarks for evaluating model performance. However, the variability in several datasets, such as ECL, is relatively homogeneous, and they do not adequately address aspects such as non-stationarity and predictability when assessing the strengths and weaknesses of models. Consequently, the development of a new benchmark is essential. Therefore, we have carefully considered factors such as domain, number of variables, frequency, non-stationarity, and predictability, and have selected a subset from the UTSD as the new benchmark. The datasets we have selected are presented in Table 3. Furthermore, we have evaluated our model along with other baseline models on these benchmarks. The results are presented in Table 4. Admittedly, relying solely on these benchmarks is not sufficient to comprehensively assess model performance. We also look forward to the proposal of more diverse and comprehensive benchmarks in the future.
Table 3: Benchmark detailed descriptions. Time Point denotes the total number of time points aggregating from all variates if multivariate. Frequency denotes the sampling interval of time points, where “-” indicates no timestamp or irregular interval. ADF Statistic denotes the Augmented Dickey-Fuller test statistics of the dataset. Forecastability denotes the forecastability of the dataset.
Domain Dataset Time Points Variates Frequency ADF Statistic Forecastability
Environment AustraliaRainfall 11.54M 3 Hourly -150.10 0.458
Transport PedestrianCounts 0.08M 1 Hourly -23.462 0.297
IoT SensorData 3.24M 18 0.002 Sec -15.892 0.917
Health BIDMC32HR 0.04M 1000 - -14.135 0.523
Table 4: Forecasting results on well-acknowledged deep forecasters and Timer, where Timer is pre-trained on the held-out datasets and then all models are superwisedly trained on the four datasets in the $672$-pred-$96$ setting.
Models Timer PatchTST iTransformer DLinear
Metric MSE MAE MSE MAE MSE MAE MSE MAE
AustraliaRainfall 0.800 0.720 0.802 0.720 0.800 0.800 0.804 0.804
PedestrianCounts 0.054 0.133 0.058 0.153 0.056 0.143 0.060 0.149
SensorData 0.049 0.094 0.056 0.094 0.052 0.091 0.057 0.111
BIDMC32HR 0.030 0.062 0.188 0.284 0.159 0.249 0.320 0.409
Domain transfer
To investigate the domain partitioning of UTSD, we use different domains of UTSD as the source and adapt the trained model to different target datasets to establish in-domain and out-of-domain transfer. The results in Table 5 indicate that in-domain transfer can further enhance the downstream performance. Additionally, as the number of downstream data samples increases, the relative improvement of pre-training will gradually diminish, and even lead to negative transfer in some out-of-domain scenarios. It provides a promising direction to develop domain-specific models.